OpenStudy (anonymous):

Can someone help me! if you help me get this right willing to give paypal money or medals!!!!!! Its a big problem!

OpenStudy (perl):

ok

OpenStudy (anonymous):

i only have 45 minutes so can you just spit out the answers.?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation. What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain. If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench.

OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. A.Use function notation that models the total price of the bracelet and how that price is based on the number of charms. Explain the reasoning behind your equation. B.What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain. C.If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. D.Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. E.As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench. Sorry here is it more clear

OpenStudy (tylerd):

ahah

OpenStudy (tylerd):

i helped someone with this same problem awhile back

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

I really need to pass this and i only have 38 minutes

OpenStudy (anonymous):

@TylerD still there?

OpenStudy (tylerd):

ya one sec

OpenStudy (anonymous):

ok

OpenStudy (tylerd):

well cant find it but

OpenStudy (anonymous):

?

OpenStudy (tylerd):

so we have f(x)=15x+85

OpenStudy (anonymous):

please man

OpenStudy (anonymous):

i got a and b already

OpenStudy (anonymous):

C i need

OpenStudy (tylerd):

ok

OpenStudy (tylerd):

whats the info for store B?

OpenStudy (anonymous):

i dont know

OpenStudy (tylerd):

well for C you make a graph so

OpenStudy (tylerd):

if f(x)=$250 this is the max money they have

OpenStudy (anonymous):

ok

OpenStudy (tylerd):

f(x)=15x+85=250

OpenStudy (tylerd):

250-85=15x where x is the number of charms

OpenStudy (tylerd):

the number of charms is your max domain value

OpenStudy (anonymous):

ok

OpenStudy (tylerd):

15x=250-85=165 165/15=x=11

OpenStudy (tylerd):

so basically make a graph from 0-250 on the y axis and 0-11 on the x axis

OpenStudy (anonymous):

here lets skip C

OpenStudy (tylerd):

f(5)=15(5)+85

OpenStudy (anonymous):

I can do it later im just in a big rush

OpenStudy (tylerd):

which is 160$

OpenStudy (anonymous):

sorry if im being rude

OpenStudy (anonymous):

Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make.

OpenStudy (tylerd):

yes, thats f(5)=160

OpenStudy (anonymous):

can you explain

OpenStudy (tylerd):

plug in 5 for x in the equation f(x)=15x+85

OpenStudy (anonymous):

ok but it says determine the cost for the bracelet so it wants us to sum it up?

OpenStudy (tylerd):

the output value is the cost

OpenStudy (tylerd):

so f(5)=cost

OpenStudy (tylerd):

we solved f(5) and got 160

OpenStudy (anonymous):

so thats the answer?

OpenStudy (tylerd):

160

OpenStudy (tylerd):

said it 4 times

OpenStudy (anonymous):

lol sorry just making sure cause this is imprtant

OpenStudy (tylerd):

now to make that graph its quite simple, can whip it up, just do from 0-11 on the X axis and 0-250 on the y axis. your line intercepts the y axis at 85 and has a slope of 15/1 rise/run

OpenStudy (anonymous):

Yea i got it already

OpenStudy (anonymous):

As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench.

OpenStudy (tylerd):

right but whats the next part of that question

OpenStudy (anonymous):

part of the graphs question?

OpenStudy (tylerd):

"As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench." isnt a math question

OpenStudy (anonymous):

Marco is wondering from which store he should buy the items so that he spends the least amount of money. Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers.

OpenStudy (anonymous):

my bad*****

OpenStudy (tylerd):

alright but i still need to know the sale for store B, what does it cost there

OpenStudy (anonymous):

160??

OpenStudy (tylerd):

no thats store A

OpenStudy (tylerd):

store a is f(x)=15x+85 we need an equation for store B

OpenStudy (anonymous):

do you know it?

OpenStudy (tylerd):

you should know it but maybe that will help

OpenStudy (anonymous):

c = 85 + 15x???

OpenStudy (anonymous):

@TylerD ?